OpenStudy (anonymous):

Given the linear equation y=-4/7x-5 find the y-coordinates of these points (-7, ),(0, ), and (7, )

OpenStudy (anonymous):

What are you confused about this question?

OpenStudy (anonymous):

k i can how to solve this problem

OpenStudy (anonymous):

i want to be sure i solved it right i got (-7,9) , (0,-5), and (7,-9)

OpenStudy (anonymous):

if x=-7 then y=-4/7(-7)-5=> y=-1

OpenStudy (anonymous):

second one is right

OpenStudy (anonymous):

third one is right.

OpenStudy (anonymous):

First one was wrong @Andresfon12 showed you how you do it

OpenStudy (anonymous):

@Romero agree

OpenStudy (anonymous):

good job mate

OpenStudy (anonymous):

ty both i will look at work to see where i went wrong

OpenStudy (anonymous):

28/7=4 4-5=-1

OpenStudy (anonymous):

ty Romero

OpenStudy (anonymous):

ty Andresfon12 i see what i did sub error